# Who is Responsible? Data, Models, or Regulations, A Comprehensive Survey on Responsible Generative AI for a Sustainable Future

### v6

@@ 4.3. Explainable AI
Recent GenAI era deployments expose failure modes (hallucination, privacy leakage, deepfakes, and tool-use errors) that cut across domains; we therefore use C1–C10 and the KPI set as a consistent lens throughout the remainder of the survey. Having established how benchmarks cover risk families, we examine explainability methods and what they can and can not do, evidence under audit. Explainable AI (XAI) comprises methods that render model behaviour and outputs intelligible to human stakeholders (Meske et al., 2022). Effective explanations serve multiple roles: they support accountability and trust, enable bias and safety analyses, and facilitate compliance with emerging governance requirements (Baker and Xiang, 2023; Arrieta et al., 2020). In practice, XAI must operate across technical, social, legal, and organisational layers rather than as a purely algorithmic add-on.
Figure 9. Post-hoc explainability within a Responsible AI (RAI) pipeline. A black-box model produces predictions $\hat{y}$ and attribution signals $I(x_{i})$, where $I(x_{i})$ denotes a feature-importance signal (e.g., gradient/saliency or a SHAP value). Interpretability pathways (feature relevance, example-based, surrogates, and visual/text explanations) feed fairness diagnostics (e.g., SPD/DI) and calibration analyses, supporting governance-aligned optimisation $\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}+\lambda\,\mathcal{L}_{\mathrm{fair}}$. For LLMs/VLMs, prompt traces and cross-modal attention provide additional observability
Scope and challenges in the GenAI
@@
XAI methods can be ante-hoc (interpretability built into the model class, e.g., sparse linear models, rule lists, decision trees) or post-hoc (explain an existing black-box locally or globally). Common post-hoc families include feature relevance (saliency/attribution, gradients, SHAP/LIME) for instance-level insight (Lundberg and Lee, 2017; Ribeiro et al., 2016a); example-based explanations (prototypes, nearest neighbours) that ground behaviour in the data manifold (Kenny et al., 2021); surrogate models that approximate decision logic with a transparent model; and visual/textual rationales (heatmaps, attention summaries) for human inspection (He et al., 2023). For LLMs and Vision-language-models (VLMs), these are complemented by prompt-engineering probes, chain-of-thought (CoT) audits, and tool-use traces (Natarajan and Nambiar, 2024). Importantly, attention maps and saliency are not proofs of causal reasoning and should be triangulated with counterfactual tests and ablations (Jain and Wallace, 2019).
@@
$\text{SPD}=P(\hat{y}=1\,|\,A=a)-P(\hat{y}=1\,|\,A=b),\quad\text{DI}=\frac{P(\hat{y}=1\,|\,A=a)}{P(\hat{y}=1\,|\,A=b)}.$
@@
$\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}(\theta)+\lambda\,\mathcal{L}_{\mathrm{fair}},$
@@
$\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\big|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\big|.$
@@
$\mathcal{E}:(x,f_{\theta},\omega)\mapsto z\in\mathcal{Z},$
@@
For foundation models, explanations should expose (i) prompt pathway elements (system/user prompts, retrieved context, tool calls), (ii) decoding state ($\omega$: temperature, top-$k/p$), and (iii) cross-modal grounding (text rationales linked to image regions or clips). These artefacts do not prove causality but enable stability and counterfactual checks and support post-hoc accountability (Rane and Paramesha, 2024; Kenny et al., 2021).
Explanations integrate naturally into risk-managed optimisation:
$\min_{\theta}\ \mathbb{E}_{x,\omega}\!\left[\ell\big(f_{\theta}(x,\omega),y\big)\right]+\lambda_{\mathrm{fair}}\mathcal{L}_{\mathrm{fair}}+\lambda_{\mathrm{xai}}\mathcal{R}\!\left(\mathcal{E}(x,f_{\theta},\omega)\right),$
@@
Limitation: Even theoretically sound explanations can mislead when they are plausible but unfaithful (e.g., over-reading attention maps), trigger Goodhart effects (Strathern, 1997b) (optimising the appearance of fairness rather than outcomes), or lack distributional validity (narratives drift across domains). Reporting stability, counterfactual checks, and provenance mitigates these risks and makes explanations suitable as audit evidence rather than marketing artefacts. Based on our synthesis findings, there is a dangerous ”Explanation Gap” in how we use these tools. While SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016b) provide useful snapshots, the literature reveals they often create ”plausible but unfaithful” stories. They give us a narrative that makes sense to a human, but because transformers are so complex and non-linear, these stories often don’t match what is actually happening under the hood (Strathern, 1997a). This creates a false sense of security that we think we’ve reached transparency when we’ve really just seen a surface-level summary. Next, we present AI-ready test-beds that computes lifecycle KPIs and logging XAI evidence as audit-ready artifacts.
@@ 4.4. AI-ready Test-beds
AI for Good (ITU) ((2017), ITU) Social good / RAI Initiative/Program UN SDG–aligned use cases; ethics and societal-benefit framing; showcases reproducible, responsible deployments.
@@
AI Commons [^3] https://ai-commons.org/ Open AI collaboration Initiative Open, transparent collaboration patterns; governance-mindful sharing; community evaluation practices (not a formal test-bed).
@@
Zeno (Zeno, 2023) AI evaluation Platform Interactive evaluation platform; enables data exploration across any data type and automatically discovers systematic failures using slice finder; supports interactive charts to compare evaluation metrics across slices and models.
Galileo (Labs, 2023) Generative AI evaluation Platform Automated evaluation metrics for GenAI (RAG and agent metrics plus safety/security metrics); real-time guardrails to block hallucinations, PII and prompt injection; captures corner cases and allows custom metrics.
Arize AI (Dong et al., 2024) AI observability Platform AI observability and LLM evaluation platform that automatically detects model issues, traces root causes via indexed training/validation/production data and fine-tunes performance; includes dashboards for evaluation/performance tracing, explainability and fairness. The Bias Tracing tool monitors fairness metrics (e.g., recall parity, false positive parity, disparate impact) and surfaces features and cohorts contributing to bias.
@@ 4.4.1. The Lifecycle Evidence Matrix: Mapping Metrics to Engineering Controls
To facilitate the implementation of our proposed evaluation infrastructure, we synthesize the relationship between KPIs, benchmarks, and technical tools. As argued throughout this survey, the effectiveness of Responsible GenAI hinges on the ability to produce portable, verifiable evidence across the lifecycle. Table 8 provides a decision-support mapping that links the KPIs defined in Section 4.2.4 to the specific testbeds and datasets that operationalize them. This matrix enables practitioners to select the appropriate ”tooling stack” based on their specific governance requirements, whether prioritizing fairness audits in Finance or robustness testing in Defense.
@@
Fairness & Bias SPD, DI, $\Delta_{\mathrm{acc}}$ SafetyBench (Zhang et al., 2024), HumaniBench (Raza et al., 2025) AIF360 (Bellamy et al., 2018), Fairlearn (Weerts et al., 2023), Arize (Dong et al., 2024)
Safety & Toxicity $E_{h}$ (High-stakes error) HarmBench (Mazeika et al., 2024), SALAD-Bench (Li et al., 2024a) Unitary (Ltd., [n. d.]), Project Moonshot (AI Verify Foundation, 2024), Galileo (Labs, 2023)
Robustness & Security $R$ (Robustness), ASR ALERT (Tedeschi et al., 2024), Rainbow Teaming (Samvelyan et al., 2024) ART (Research, [n. d.]), PromptLayer (PromptLayer, 2022), ToolSandbox (Lu et al., 2024a)
Privacy $P$ (Compliance rate) PrivLM-Bench (Li et al., 2024b), PrivacyLens (Shao and et al., 2024) Privacy Meter (Privacy and Lab, [n. d.]), Google DP (Google, [n. d.]), SecretFlow (Group, [n. d.])
Explainability $E(x)$, LFG (Faithfulness) HELM (Liang et al., 2023), XSTest (Röttger et al., 2024) LIT (Google PAIR, [n. d.]), InterpretML, AIX360 (Arya et al., 2019)
Sustainability $E_{\mathrm{kWh}}$, $C_{\mathrm{CO_{2}e}}$ CodeCarbon (Song et al., 2025) Evidently (Evidently AI, [n. d.]), Arize (Dong et al., 2024), Zeno (Zeno, 2023)
@@ 4.4.2. Turning Metrics into Assurance: The ResGenAI Audit Loop
The technical efficacy of ResGenAI remains localized unless it can be translated into portable audit evidence. As visualized in Fig. 10, our framework proposes a three layer loop that bridges the Evaluation Gap identified in Section 6.4. In this cycle, the technical KPIs defined in Section 4.2.4 act as the raw sensory data for the ecosystem. These metrics are packaged into standardized audit artifacts such as model cards and red teaming logs which serve as the portable evidence required by compliance auditors under the EU AI Act (European Union, 2024) or NIST AI RMF (of Standards and, NIST). By establishing this formal “chain of trust”, organizations can move beyond static safety checks toward continuous stewardship where regulatory feedback directly informs the recalibration of technical metrics.
@@ 5. Responsible Generative AI Application Across Domains
GenAI systems are increasingly embedded in decision-making processes across diverse domains such as healthcare, finance, human resource, defense and education. While these systems offer significant utility, they also introduce new categories of risks and challenges, as outlined earlier. The challenges and mitigation strategies synthesized in this review are broadly relevant, but their salience and priority vary substantially by sectoral context. We translate the above findings into stakeholder-aligned practices and a deployment-oriented roadmap, with examples and decision pointers. Each subsection below examines domain-specific applications, challenges, and strategies through the lens of our governance-aligned rubric (C1–C10, Section 4.2.1) and lifecycle KPIs (Section 4.2.4). For every domain we (i) identify which rubric criteria are most salient, (ii) assess how the benchmarks in Tables 46 cover those risks, and (iii) indicate which KPIs practitioners should prioritize. Table 9 consolidates this analysis, mapping each sector to its primary criteria gaps and priority KPIs.
@@ 5.1. Healthcare
Applications
Generative models improve efficiency in clinical note generation and triage chatbots (Alagha and Ahn, 2023), they raise concerns about accuracy, liability, and bias in medical decision support. Disease detection and drug discovery pipelines (Liu et al., 2023c) illustrate GenAI’s potential for accelerating science, yet demand rigorous validation, explainability, and regulatory alignment before clinical deployment. Synthetic health data helps mitigate privacy risks and dataset imbalance (Walsh et al., 2023), but governance frameworks must ensure that synthetic outputs do not inadvertently reinforce hidden biases. Similarly, tools like Watson for Oncology highlight the need for transparency and interpretability in high-stakes decision-making (Memorial Sloan Kettering Cancer Center, 2024). Even promising breakthroughs, such as GenAI-designed therapeutics, require human oversight on trial integrity, equitable access, and long-term monitoring. In this way, healthcare exemplifies both the promise of GenAI and the critical role of RAI practices, balancing innovation with safety, accountability, and fairness.
ResGenAI challenges
@@
Responsible adoption of GenAI in healthcare is constrained by privacy violations (Shiraz and Lee, 2019), hallucination of critical clinical facts (Li et al., 2023b), algorithmic bias in diagnostics (Obermeyer et al., 2019), unclear accountability for errors (Kaminski, 2022), opacity of generative models (Ghorbani et al., 2019), and difficulty of compliance with traditional regulations (Price and Gerke, 2022). Empirical studies highlight these risks, for example, a 2023 evaluation of 450 consultation transcripts using GPT-4 reported a 1.47% hallucination rate, with 44% of errors classified as major and concentrated in treatment plans (Li et al., 2023b); Rosbach et al. further found a 7% automation-bias error rate in a pathology study where clinicians overturned correct diagnoses by over-trusting AI outputs (Rosbach et al., 2024); and a systematic review of 778 studies showed that explainable AI can both increase and erode clinician trust, underscoring regulatory and accountability challenges (Tjoa and Guan, 2023).
Evaluation context
Mapping these risks to our rubric, the dominant healthcare concerns correspond to C4 (hallucination of clinical facts), C1 (diagnostic bias), and C6 (patient data leakage). Only 7 of 21 benchmarks in Table 4 score above zero on C4 (Misinformation/ Disinformation), and none specifically target medical hallucination. For privacy, only PrivLM-Bench and PrivacyLens provide strong C6 (Privacy) coverage, however, neither addresses HIPAA-specific leakage vectors such as clinical note memorization. C7 (System-level failure) is critical for RAG-based clinical pipelines and is tested by only 4 of 21 benchmarks. The KPIs most relevant to this domain are the high-stakes error rate $E_{h}$ (KPI 9), operationalized as hallucinated clinical facts per $N$ consultations, data privacy compliance $P$ (KPI 2) under HIPAA, and fairness metrics SPD/DI (KPIs 3–4) stratified by patient demographics.
ResGenAI strategies
Several strategies are being employed to enable responsible GenAI adoption in healthcare. Grounding methods (C4, KPI 9) such as RAG can help reduce hallucinations by linking model outputs to verifiable medical sources. Effectiveness should be tracked via $E_{h}$; a reported 1.47% hallucination rate provides a baseline. Human-in-the-loop (HITL) validation (C7, C8, KPI 10) integrates clinician oversight into diagnostic and documentation workflows to mitigate automation bias. Audit frequency $F_{a}$ and resolution time $T_{r}$ should be reported.Privacy-preserving approaches (C6, KPI 2), including differential privacy and federated learning, protect sensitive health records while supporting multi-institutional collaboration. Fairness-aware data augmentation and structured fairness audits (C1, KPIs 3–5) are being applied to improve equity in clinical outcomes. HumaniBench (C1=2 in Table 4) offers a starting template but lacks clinical subgroup stratification. Post-hoc explainability (C9, KPIs 7–8) techniques such as SHAP and LIME are critical for clinician trust.Explanation quality should be assessed via the local faithfulness gap LFG (KPI 8). Alignment with regulatory frameworks (C10) like FDA guidance ensures safe and accountable deployment. Table 6 shows that only HELM and MLCommons v0.5 provide partial transparency support relevant to FDA conformity assessment.

### v7

@@ 4.3. Explainable AI
Explainable AI (XAI) comprises methods that render model behaviour and outputs intelligible to human stakeholders (Meske et al., 2022). Effective explanations serve multiple roles: they support accountability and trust, enable bias and safety analyses, and facilitate compliance with emerging governance requirements. In practice, XAI must operate across technical, social, legal, and organisational layers rather than as a purely algorithmic add-on.
Figure 8. Post-hoc explainability within a ResGenAI. A black-box model produces predictions $\hat{y}$ and attribution signals $I(x_{i})$, where $I(x_{i})$ denotes a feature-importance signal (e.g., gradient/saliency or a SHAP value). Interpretability pathways (feature relevance, example-based, surrogates, and visual/text explanations) feed fairness diagnostics (e.g., SPD/DI) and calibration analyses, supporting governance-aligned optimisation $\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}+\lambda\,\mathcal{L}_{\mathrm{fair}}$. For LLMs/VLMs, prompt traces and cross-modal attention provide observability
@@
XAI methods can be ante-hoc (interpretability built into the model class, e.g., sparse linear models, rule lists, decision trees) or post-hoc (explain an existing black-box locally or globally). Common post-hoc families include feature relevance (saliency/attribution, gradients, SHAP/LIME) for instance-level insight (Lundberg and Lee, 2017; Ribeiro et al., 2016b); example-based explanations (prototypes, nearest neighbours) that ground behaviour in the data manifold (Kenny et al., 2021); surrogate models that approximate decision logic with a transparent model; and visual/textual rationales (heatmaps, attention summaries) for human inspection (He et al., 2023). For LLMs and VLMs, these are complemented by prompt-engineering probes, chain-of-thought (CoT) audits, and tool-use traces (Natarajan and Nambiar, 2024). Importantly, attention maps and saliency are not proofs of causal reasoning and should be triangulated with counterfactual tests and ablations (Jain and Wallace, 2019).
@@
(14) $\text{SPD}=P(\hat{y}=1\,|\,A=a)-P(\hat{y}=1\,|\,A=b),\quad\text{DI}=\frac{P(\hat{y}=1\,|\,A=a)}{P(\hat{y}=1\,|\,A=b)}.$
@@
(15) $\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}(\theta)+\lambda\,\mathcal{L}_{\mathrm{fair}},$
@@
(16) $\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\big|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\big|.$
@@
(17) $\mathcal{E}:(x,f_{\theta},\omega)\mapsto z\in\mathcal{Z},$
@@
For foundation models, explanations should expose (i) prompt pathway elements (system/user prompts, retrieved context, tool calls), (ii) decoding state ($\omega$: temperature, top-$k/p$), and (iii) cross-modal grounding (text rationales linked to image regions or clips). These artefacts do not prove causality but enable stability and counterfactual checks and support post-hoc accountability. Explanations integrate naturally into risk-managed optimisation:
(18) $\min_{\theta}\ \mathbb{E}_{x,\omega}\!\left[\ell\big(f_{\theta}(x,\omega),y\big)\right]+\lambda_{\mathrm{fair}}\mathcal{L}_{\mathrm{fair}}+\lambda_{\mathrm{xai}}\mathcal{R}\!\left(\mathcal{E}(x,f_{\theta},\omega)\right),$
@@
Limitation: Even theoretically sound explanations can mislead when they are plausible but unfaithful (e.g., over-reading attention maps), trigger Goodhart effects (Strathern, 1997) (optimising the appearance of fairness rather than outcomes), or lack distributional validity (narratives drift across domains). Reporting stability, counterfactual checks, and provenance mitigates these risks and makes explanations suitable as audit evidence rather than marketing artefacts. Based on our synthesis findings, there is a dangerous ”Explanation Gap” in how we use these tools. While SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016a) provide useful snapshots, the literature reveals they often create ”plausible but unfaithful” stories. They give us a narrative that makes sense to a human, but because transformers are so complex and non-linear, these stories often don’t match what is actually happening under the hood (Strathern, 1997). This creates a false sense of security that we think we’ve reached transparency when we’ve really just seen a surface-level summary. Next, we present AI-ready test-beds that computes lifecycle KPIs and logging XAI evidence as audit-ready artifacts.
@@ 4.4. AI-ready Test-beds
AI for Good (ITU) ((ITU), 2017) Social good / RAI Initiative/Program UN SDG–aligned use cases; ethics and societal-benefit framing; showcases reproducible, responsible deployments.
@@
AI Commons Open AI collaboration Initiative Open, transparent collaboration patterns; governance-mindful sharing; community evaluation practices (not a formal test-bed).
@@ 4.4.1. The Lifecycle Evidence Matrix: Mapping Metrics to Engineering Controls
To facilitate the implementation of our proposed evaluation, we synthesize the relationship between KPIs, benchmarks, and technical tools. Table. Table 7 provides a decision-support mapping that links the KPIs defined in Section 4.2.4 to the specific testbeds and datasets that operationalize them. This matrix enables practitioners to select the appropriate ”tooling stack” based on their specific governance requirements, whether prioritizing fairness audits in Finance or robustness testing in Defense.
@@
Fairness & Bias SPD, DI, $\Delta_{\mathrm{acc}}$ SafetyBench (Zhang et al., 2024), HumaniBench (Raza et al., 2025) AIF360 (Bellamy et al., 2018), Fairlearn (Weerts et al., 2023)
Safety & Toxicity $E_{h}$ (High-stakes error) HarmBench (Mazeika et al., 2024), SALAD-Bench (Li et al., 2024b) Unitary (Ltd., ), Project Moonshot (AI Verify Foundation, 2024), Galileo (Labs, 2023)
Robustness & Security $R$ (Robustness), ASR ALERT (Tedeschi et al., 2024), Rainbow Teaming (Samvelyan et al., 2024) ART (Research, ), PromptLayer (PromptLayer, 2022), ToolSandbox (Lu et al., 2024a)
Privacy $P$ (Compliance rate) PrivLM-Bench (Li et al., 2024a), PrivacyLens (Shao and et al., 2024) Privacy Meter (Privacy and Lab, ), Google DP (Google, 2025), SecretFlow (Group, )
Explainability $E(x)$, LFG (Faithfulness) HELM (Liang et al., 2023), XSTest (Röttger et al., 2024) LIT (Google PAIR, ), InterpretML, AIX360 (Arya et al., 2019)
Sustainability $E_{\mathrm{kWh}}$, $C_{\mathrm{CO_{2}e}}$ CodeCarbon (Song et al., 2025) Evidently (Evidently AI, ),
@@ 4.4.2. Turning Metrics into Assurance: The ResGenAI Audit Loop
The technical efficacy of ResGenAI remains localized unless it can be translated into portable audit evidence. As visualized in Fig. 9, our framework proposes a three layer loop that bridges the ”Evaluation Gap identified”. In this cycle, the technical KPIs defined in Section 4.2.4 act as the raw sensory data for the ecosystem. These metrics are packaged into standardized audit artifacts such as model cards and red teaming logs which serve as the portable evidence required by compliance auditors under the EU AI Act or NIST AI RMF. By establishing this formal “chain of trust”, organizations can move beyond static safety checks toward continuous stewardship where regulatory feedback directly informs the recalibration of technical metrics.
@@
Next, we discuss the ResGenAI across domains and map how these aforementioned methods can be applied in these applications.
@@ 5. Responsible Generative AI Application Across Domains
GenAI systems are increasingly embedded in decision-making processes across diverse domains such as healthcare, finance, education, defense, arts, and agriculture. Each subsection below examines domain-specific risks and strategies. For every domain we identify which rubric criteria (C1–C10) are most salient and indicate which KPIs practitioners should prioritize. Appendix Table 6 consolidates this analysis, mapping each sector to its primary criteria gaps and priority KPIs.
Cross-cutting strategies.
Several responsible deployment strategies recur across nearly all high-stakes domains and form a shared baseline. Fairness audits (C1, KPIs 3–5) using SPD, DI, and accuracy-parity metrics should be conducted at regular intervals, stratified by relevant demographic or contextual variables. Privacy-preserving techniques (C6, KPI 2), including differential privacy, federated learning, and secure multiparty computation, enable cross-institutional collaboration without exposing sensitive records. Post-hoc explainability (C9, KPIs 7–8) via SHAP, LIME, or chain-of-thought audits supports stakeholder trust and regulatory interrogation of opaque models. Human-in-the-loop (HITL) oversight (C7, KPI 10) embeds expert review into high-consequence decision points, with audit frequency $F_{a}$ and resolution time $T_{r}$ tracked as governance evidence. Regulatory alignment (C10) anchors deployments within enforceable legal standards (e.g., EU AI Act, NIST AI RMF, sector-specific mandates). The subsections below focus on what is distinctive about each domain’s risk profile and mitigation needs beyond this shared baseline.
@@ 5.1. Healthcare
Applications and challenges. Generative models improve efficiency in clinical note generation and triage chatbots, yet raise concerns about accuracy, liability, and bias in medical decision support. Responsible adoption is constrained by hallucination of critical clinical facts, algorithmic bias in diagnostics, privacy violations, unclear accountability, and difficulty of compliance with sector-specific regulations. A 2023 evaluation of 450 consultation transcripts using GPT-4 reported a 1.47% hallucination rate, with 44% of errors classified as major and concentrated in treatment plans (Li et al., 2023a). A systematic review of 778 studies also showed that explainable AI can both increase and erode clinician trust, underscoring regulatory and accountability challenges (Tjoa and Guan, 2023; Raza et al., 2023).
Domain-specific strategies. Beyond the shared baseline, healthcare demands grounding via retrieval-augmented generation (RAG) to link outputs to verifiable medical sources, with effectiveness tracked through the high-stakes error rate $E_{h}$ (KPI 9) operationalized as hallucinated clinical facts per $N$ consultations. Fairness-aware data augmentation stratified by patient demographics addresses diagnostic disparities that generic audits may miss. HITL validation is particularly critical given documented automation-bias effects where clinicians overturn correct diagnoses by over-trusting AI outputs. Alignment with FDA guidance and HIPAA-specific privacy requirements (e.g., clinical-note memorization vectors) distinguishes this domain from the general privacy baseline.